OpenStudy (jenniferjuice):

jesus someone please help Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. y = a(x-h)² + k a= 4 -(-4) h=8 k=16 y=4(x-8)²+16 Using complete sentences, explain how to convert your birthday function into standard form. 8/16/99

jimthompson5910 (jim_thompson5910):

This question makes no sense at all (no offense)

jimthompson5910 (jim_thompson5910):

but I'm guessing all you do is plug in your birthday into y = a(x-h)² + k and that's all they want? Not sure

OpenStudy (jenniferjuice):

i know it doesnt make sense >.<

jimthompson5910 (jim_thompson5910):

I'd just skip that problem and move onto other hw problems. Idk what else to say

jimthompson5910 (jim_thompson5910):

ok what values of a, h, k are you using?

OpenStudy (jenniferjuice):

8 16 and 99 im assuming

jimthompson5910 (jim_thompson5910):

oh wait, it says "a" has to be between -4 and 4

jimthompson5910 (jim_thompson5910):

so pick another "a" value

OpenStudy (jenniferjuice):

4?

jimthompson5910 (jim_thompson5910):

that works

jimthompson5910 (jim_thompson5910):

a = 4 h = 16 k = 99 so... \[\Large y = a(x-h)^2 + k\] \[\Large y = 4(x-16)^2 + 99\]

jimthompson5910 (jim_thompson5910):

which takes care of part a)

jimthompson5910 (jim_thompson5910):

Then we convert to standard form (to answer part b) \[\Large y = 4(x-16)^2 + 99\] \[\Large y = 4(x-16)(x-16) + 99\] \[\Large y = 4(x^2-32x+256) + 99\] \[\Large y = 4x^2-128x+1024 + 99\] \[\Large y = 4x^2-128x+1123\]

OpenStudy (jenniferjuice):

oh my gosh thank you so much for your help

jimthompson5910 (jim_thompson5910):

np

OpenStudy (jenniferjuice):

1. a = 4 h = 16 k = 99 y=a(x-h)^2+k y=4(x-16)^2+99 2. y=4(x-16)^2+99 y=4(x-16)(x-16)+99 y=4

OpenStudy (jenniferjuice):

1. a = 4 h = 16 k = 99 y=a(x-h)^2+k y=4(x-16)^2+99 2. y=4(x-16)^2+99 y=4(x-16)(x-16)+99 y=4(x^2 – 32x+ 256)+99 y=4(x^2-